# Characterisation of the Bedretto Underground Site for Fundamental Physics Experiments

### v1

@@ 1 Introduction
Underground laboratories provide uniquely background-free environments and are indispensable tools to understand fundamental physics processes. The massive rock overburden shields effectively even from the highest energetic cosmic rays, they are seismically very quiet since most vibrations travel at the surface and are attenuated quickly at depth, and Newtonian noise is reduced by the very large mass of rock surrounding the laboratory. Examples of underground experiments include rare-event searches such as the search for dark matter and neutrinoless double beta decay, rare-decay nuclear physics measurements, low-background material assay, environmental monitoring, and optimal sensitivity and stability of quantum sensors. Furthermore, the stable, low-background, and low-vibration environment offered by underground locations is highly beneficial for gravitational wave detection. Several major deep underground laboratories exist but none in Switzerland. These facilities often have to prioritise large-scale experiments and offer only limited space for R&D setups and prototyping. Furthermore, their often remote locations and necessary travel leads to logistical and practical overhead. The Bedretto site in the Canton of Ticino, Switzerland, provides an ideal location to establish a Swiss deep underground laboratory for fundamental physics. As one of the deepest and quietest laboratories in Europe, it is easily accessible from all Swiss research institutions and features horizontal access to the underground site.
The Bedretto tunnel is an annex to the Furka railway starting in Bedretto, Ticino, Switzerland. The tunnel was used to remove excavation material during construction and as an evacuation route of the railway.
@@
Figure 1: The Bedretto tunnel and the profile of the mountain with the different rock formations overlaid rast2022geology; nasa2019aster. The black markers indicate the gamma-ray rate as measured using a 3 in NaI(Tl) detector. The red line indicates the site of the planned fundamental physics lab. Serendipitously, the laboratory’s intended location the gamma-ray flux is close to minimal while the overburden is maximal. The locations of rock niches are indicated in gray in the bottom panel.
To determine the suitability of the site for low background physics, a comprehensive site characterisation campaign was performed in summer and autumn 2025. We conducted measurements of key background contributions such as radiative backgrounds, specifically cosmic muon, neutron, and $\gamma$-ray fluxes, and radon activity, as well as magnetic and seismic background levels.
@@ 2 Atmospheric Muon Flux
Muons and neutrons from muon-induced spallation processes constitute a major and difficult to shield background for rare event searches. As minimum-ionizing particles, high-energy atmospheric muons are able to penetrate large amounts of matter and only underground laboratories allow for significant reduction of the cosmic muon flux. Underground laboratories are typically characterised in terms of metre-water-equivalent (m.w.e) in their ability to attenuate the cosmic muon flux Heise:2022iaf; Bruno:2019vad; Woodley:2024eln; LSM_Muon, see Fig. 2.
@@ 2.1 Muon Telescope
The atmospheric muon flux was measured at the TM3500 site utilizing a telescope composed of four EJ-200 plastic scintillator panels EJ-200, each with an area of 625 cm2 and 3.8 cm in thickness. An array of four Onsemi C-series silicon photomultipliers (SiPM) was attached to each panel, biased with a CAEN SiPM power supply at 31 V CAENsipm and read out with a CAEN DT5730 500 MHz digitizer employing DPP-PHA firmware CAEN5730; CAEN-DPP-PHA. Lead plates, about $2$ cm thick, were placed between the plastic scintillator to minimise spurious gamma and Compton scattering backgrounds.
@@ 3 Gamma Ray Flux
To asses the natural radioactivity of the granite wall and inform Monte-Carlo-based predictions of gamma and neutron fluxes, we measured different granite samples with the Gator high purity germanium detector (HPGe) facility at LNGS.
@@ 8 Discussion
The sites allows for easy and fast excavations and expansion of existing underground space. Our plan is to developed a sizeable O(100 m2) underground clean laboratory for underground R&D for rare event searches, quantum sensing, gravitational waves and others quickly. Future expansion are possible including excavation of O(100m-1km) tunnels and shafts for, e.g., atom interferometry, gravitational wave and neutrino detectors. The site offers excellent accessibility to leading research institutions and major cities such as Zürich, Geneva, and Milano. Additional logistical benefits are the horizontal access, the capacity to transport up to 18 t of mass at once and existing surface infrastructure.

### v2

@@ 1 Introduction
Underground laboratories provide uniquely background-free environments and are indispensable tools to understand fundamental physics processes. The massive rock overburden shields effectively even from the highest energetic cosmic rays, they are seismically very quiet since most vibrations travel at the surface and are attenuated quickly at depth, and Newtonian noise is reduced by the very large mass of rock surrounding the laboratory. Examples of underground experiments include rare-event searches such as the search for dark matter and neutrinoless double beta decay, rare-decay nuclear physics measurements, low-background material assay, environmental monitoring, and optimal sensitivity and stability of quantum sensors. Furthermore, the stable, low-background, and low-vibration environment offered by underground locations is highly beneficial for gravitational wave detection. Several major deep underground laboratories exist, but none in Switzerland. These facilities often have to prioritise large-scale experiments and offer only limited space for R&D setups and prototyping. Furthermore, their often remote locations and necessary travel lead to logistical and practical overhead. The Bedretto site in the Canton of Ticino, Switzerland, provides an ideal location to establish a Swiss deep underground laboratory for fundamental physics. As one of the deepest and quietest laboratories in Europe, it is easily accessible from all Swiss research institutions and features horizontal access to the underground site.
The Bedretto tunnel is an annex to the Furka railway starting in Bedretto, Ticino, Switzerland. The tunnel was used to remove excavation material during construction and as an evacuation route for the railway.
@@
Figure 1: The Bedretto tunnel and the profile of the mountain with the different rock formations overlaid rast2022geology; nasa2019aster. The black markers indicate the gamma-ray rate as measured using a 3 in NaI(Tl) detector. The red line indicates the site of the planned fundamental physics lab. Serendipitously, the laboratory’s intended location, the gamma-ray flux is close to minimal while the overburden is maximal. The locations of rock niches are indicated in grey in the bottom panel.
To determine the suitability of the site for low-background physics, a comprehensive site characterisation campaign was performed in summer and autumn 2025. We conducted measurements of key background contributions such as radiative backgrounds, specifically cosmic muon, neutron, and $\gamma$-ray fluxes, and radon activity, as well as magnetic and seismic background levels.
@@ 2 Atmospheric Muon Flux
Muons and neutrons from muon-induced spallation processes constitute a major and difficult to shield background for rare event searches. As minimum-ionising particles, high-energy atmospheric muons are able to penetrate large amounts of matter, and only underground laboratories allow for a significant reduction of the cosmic muon flux. Underground laboratories are typically characterised in terms of metre-water-equivalent (m.w.e) in their ability to attenuate the cosmic muon flux Heise:2022iaf; Bruno:2019vad; Woodley:2024eln; LSM_Muon, see Fig. 2.
@@ 2.1 Muon Telescope
The atmospheric muon flux was measured at the TM3500 site utilising a telescope composed of four EJ-200 plastic scintillator panels EJ-200, each with an area of 625 cm2 and 3.8 cm in thickness. An array of four Onsemi C-series silicon photomultipliers (SiPM) was attached to each panel, biased with a CAEN SiPM power supply at 31 V CAENsipm and read out with a CAEN DT5730 500 MHz digitizer employing DPP-PHA firmware CAEN5730; CAEN-DPP-PHA. Lead plates, about $2$ cm thick, were placed between the plastic scintillator to minimise spurious gamma and Compton scattering backgrounds.
@@ 3 Gamma Ray Flux
To assess the natural radioactivity of the granite wall and inform Monte-Carlo-based predictions of gamma and neutron fluxes, we measured different granite samples with the Gator high-purity germanium detector (HPGe) facility at LNGS.
@@ 8 Discussion
The site allows for easy and fast excavations and expansion of existing underground space. Our plan is to developed a sizeable O(100 m2) underground clean laboratory for underground R&D for rare event searches, quantum sensing, gravitational waves and others quickly. Future expansion are possible including excavation of O(100m-1km) tunnels and shafts for, e.g., atom interferometry, gravitational wave and neutrino detectors. The site offers excellent accessibility to leading research institutions and major cities such as Zürich, Geneva, and Milano. Additional logistical benefits are the horizontal access, the capacity to transport up to 18 t of mass at once and existing surface infrastructure.